OpenStudy (anonymous):

I WILL MEDAL!!! A is the midpoint of XY Find: XA Find: AY and XY X--^3x---A---^5x-6--Y

OpenStudy (chrisisdope1):

AY = 2(5x-6), XY = XA + AY

OpenStudy (anonymous):

but whats XA?

OpenStudy (chrisisdope1):

XA In Your Problem!!!

OpenStudy (anonymous):

no XA has to equal an exact whole number.....